Consider the function f(x)=(3x+6)/(3x+2). For this function there are two important intervals: (−∞,A) and (A,∞) where the function is not defined at A. Determine A.

A=

For each of the following intervals, tell whether f(x) is increasing (input INC ) or decreasing (type in DEC ).

(−∞,A):
(A,∞):

Note that this function has no inflection points, but we can still consider its concavity. For each of the following intervals, tell whether f(x) is concave up (input CU ) or concave down (type in CD ).

(−∞,A):
(A,∞):
1
Expert's answer
2012-11-05T09:53:26-0500
we find A from denominator cant be zero so
3x+2=0 so
x=-3/2 A=-3/2
f'=-12/(3x+2)<0
x not=A so f DEC on the both (−∞,A), (A,∞):
f"=72/(3x+2)^3x belongs (−∞,A) f"<0 then CU
x belongs (A,∞) f">0 then CD